4. Experiments show that the actual number of muons that reach the base of Mt. Washington is 400.




a. According to your muon clock graph, how much time has elapsed if 400 muons reach the base of Mt. Washington?




b. By what factor do the times you found differ?




c. Suggest an explanation for this difference.

1
Expert's answer
2022-03-20T18:47:35-0400

"t=\frac sv=6.73~\mu s,"

"N=N_0(1-2^{-\frac tT})=475."